SolveIt is under development
SolveItNCERT · CBSE · NEET

NCERT Exemplar · Class 12 Mathematics Probability

108 questions · 108 still being checked

EXERCISE 13.3 61–70 (part 7 of 11)

  1. Choose the correct answer from the given four options in each of the exercises from $\displaystyle 56$ to 82.

    Exercise 61

    If P(A)=0.4,P(B)=0.8\displaystyle \mathrm{P}(\mathrm{A})=0.4, \mathrm{P}(\mathrm{B})=0.8 and P(B∣A)=0.6\displaystyle \mathrm{P}(\mathrm{B} \mid \mathrm{A})=0.6, then P(A∪B)\displaystyle \mathrm{P}(\mathrm{A} \cup \mathrm{B}) is equal to
    (A)
    0.24\displaystyle 24
    (B)
    0.3\displaystyle 3
    (C)
    0.48\displaystyle 48
    (D)
    0.96\displaystyle 96

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    D
    (D) \(\displaystyle 0.96\)\[P(A\cap B)=P(B\mid A)\,P(A)=0.6\times0.4=0.24 \] \[P(A\cup B)=P(A)+P(B)-P(A\cap B) \] \[=0.4+0.8-0.24=0.96 \]
  2. Exercise 62

    If A and B are two events and A≠ϕ,B≠ϕ\displaystyle \mathrm{A} \neq \phi, \mathrm{B} \neq \phi, then
    (A)
    P(A∣B)=P(A)⋅P(B)\displaystyle \mathrm{P}(\mathrm{A} \mid \mathrm{B})=\mathrm{P}(\mathrm{A}) \cdot \mathrm{P}(\mathrm{B})
    (B)
    P(A∣B)=P(A∩B)P(B)\displaystyle \mathrm{P}(\mathrm{A} \mid \mathrm{B})=\frac{\mathrm{P}(\mathrm{A} \cap \mathrm{B})}{\mathrm{P}(\mathrm{B})}
    (C)
    P(A∣B)⋅P(B∣A)=1\displaystyle \mathrm{P}(\mathrm{A} \mid \mathrm{B}) \cdot \mathrm{P}(\mathrm{B} \mid \mathrm{A})=1
    (D)
    P(A∣B)=P(A)∣P(B)\displaystyle \mathrm{P}(\mathrm{A} \mid \mathrm{B})=\mathrm{P}(\mathrm{A}) \mid \mathrm{P}(\mathrm{B})

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    B
    (B) This is the definition of conditional probability.\[P(A\mid B)=\frac{P(A\cap B)}{P(B)} \]For disjoint events of positive probability, \(\displaystyle P(A\mid B)=0\), so (A) and (C) fail; (D) is not a probability statement.
  3. Exercise 63

    A and B are events such that P(A)=0.4,P(B)=0.3\displaystyle \mathrm{P}(\mathrm{A})=0.4, \mathrm{P}(\mathrm{B})=0.3 and P(A∪B)=0.5\displaystyle \mathrm{P}(\mathrm{A} \cup \mathrm{B})=0.5. Then P(B′∩A)\displaystyle \mathrm{P}\left(\mathrm{B}^{\prime} \cap \mathrm{A}\right) equals
    (A)
    23\displaystyle \frac{2}{3}
    (B)
    12\displaystyle \frac{1}{2}
    (C)
    310\displaystyle \frac{3}{10}
    (D)
    15\displaystyle \frac{1}{5}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    D
    (D) \(\displaystyle \dfrac{1}{5}\)\[P(A\cap B)=P(A)+P(B)-P(A\cup B) \] \[=0.4+0.3-0.5=0.2 \] \[P(B'\cap A)=P(A)-P(A\cap B) \] \[=0.4-0.2=0.2=\frac{1}{5} \]
  4. Exercise 64

    You are given that A and B are two events such that P(B)=35,P(A∣B)=12\displaystyle \mathrm{P}(\mathrm{B})=\frac{3}{5}, \mathrm{P}(\mathrm{A} \mid \mathrm{B})=\frac{1}{2} and P(A∪B)=45\displaystyle \mathrm{P}(\mathrm{A} \cup \mathrm{B})=\frac{4}{5}, then P(A)\displaystyle \mathrm{P}(\mathrm{A}) equals
    (A)
    310\displaystyle \frac{3}{10}
    (B)
    15\displaystyle \frac{1}{5}
    (C)
    12\displaystyle \frac{1}{2}
    (D)
    35\displaystyle \frac{3}{5}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    C
    (C) \(\displaystyle \dfrac{1}{2}\)\[P(A\cap B)=P(A\mid B)\,P(B)=\frac{1}{2}\cdot\frac{3}{5}=\frac{3}{10} \] \[P(A\cup B)=P(A)+P(B)-P(A\cap B) \] \[\frac{4}{5}=P(A)+\frac{3}{5}-\frac{3}{10} \] \[P(A)=\frac{4}{5}-\frac{3}{5}+\frac{3}{10}=\frac{1}{2} \]
  5. Exercise 65

    In Exercise 64\displaystyle 64 above, P(B∣A′)\displaystyle \mathrm{P}\left(\mathrm{B} \mid \mathrm{A}^{\prime}\right) is equal to
    (A)
    15\displaystyle \frac{1}{5}
    (B)
    310\displaystyle \frac{3}{10}
    (C)
    12\displaystyle \frac{1}{2}
    (D)
    35\displaystyle \frac{3}{5}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    D
    (D) \(\displaystyle \dfrac{3}{5}\)\[P(A\cap B)=P(A\mid B)\,P(B)=\frac12\cdot\frac35=\frac{3}{10} \] \[P(A)=P(A\cup B)-P(B)+P(A\cap B)=\frac45-\frac35+\frac{3}{10}=\frac12 \] \[P(A'\cap B)=P(B)-P(A\cap B)=\frac35-\frac{3}{10}=\frac{3}{10} \] \[P(B\mid A')=\frac{P(A'\cap B)}{P(A')}=\frac{3/10}{1-\frac12}=\frac35 \]
  6. Exercise 66

    If P(B)=35,P(A∣B)=12\displaystyle \mathrm{P}(\mathrm{B})=\frac{3}{5}, \mathrm{P}(\mathrm{A} \mid \mathrm{B})=\frac{1}{2} and P(A∪B)=45\displaystyle \mathrm{P}(\mathrm{A} \cup \mathrm{B})=\frac{4}{5}, then P(A∪B)′+P(A′∪B)=\displaystyle \mathrm{P}(\mathrm{A} \cup \mathrm{B})^{\prime}+\mathrm{P}\left(\mathrm{A}^{\prime} \cup \mathrm{B}\right)=
    (A)
    15\displaystyle \frac{1}{5}
    (B)
    45\displaystyle \frac{4}{5}
    (C)
    12\displaystyle \frac{1}{2}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    D
    (D) \(\displaystyle 1\)\[P(A\cap B)=P(A\mid B)\,P(B)=\frac12\cdot\frac35=\frac{3}{10} \] \[P(A)=P(A\cup B)-P(B)+P(A\cap B)=\frac45-\frac35+\frac{3}{10}=\frac12 \] \[P\big((A\cup B)'\big)=1-\frac45=\frac15 \] \[P(A'\cup B)=P(A')+P(B)-P(A'\cap B)=\frac12+\frac35-\frac{3}{10}=\frac45 \] \[P\big((A\cup B)'\big)+P(A'\cup B)=\frac15+\frac45=1 \]
  7. Exercise 67

    Let P(A)=713,P(B)=913\displaystyle \mathrm{P}(\mathrm{A})=\frac{7}{13}, \mathrm{P}(\mathrm{B})=\frac{9}{13} and P(A∩B)=413\displaystyle \mathrm{P}(\mathrm{A} \cap \mathrm{B})=\frac{4}{13}. Then P(A′∣B)\displaystyle \mathrm{P}\left(\mathrm{A}^{\prime} \mid \mathrm{B}\right) is equal to
    (A)
    613\displaystyle \frac{6}{13}
    (B)
    413\displaystyle \frac{4}{13}
    (C)
    49\displaystyle \frac{4}{9}
    (D)
    59\displaystyle \frac{5}{9}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    D
    (D) \(\displaystyle \dfrac{5}{9}\)\[P(A'\cap B)=P(B)-P(A\cap B)=\frac{9}{13}-\frac{4}{13}=\frac{5}{13} \] \[P(A'\mid B)=\frac{P(A'\cap B)}{P(B)}=\frac{5/13}{9/13}=\frac59 \]
  8. Exercise 68

    If A and B are such events that P(A)>0\displaystyle \mathrm{P}(\mathrm{A})>0 and P(B)≠1\displaystyle \mathrm{P}(\mathrm{B}) \neq 1, then P(A′∣B′)\displaystyle \mathrm{P}\left(\mathrm{A}^{\prime} \mid \mathrm{B}^{\prime}\right) equals.
    (A)
    1−P(A∣B)\displaystyle 1-\mathrm{P}(\mathrm{A} \mid \mathrm{B})
    (B)
    1−P(A′∣B)\displaystyle 1-\mathrm{P}\left(\mathrm{A}^{\prime} \mid \mathrm{B}\right)
    (C)
    1−P(A∪B)P(B′)\displaystyle \frac{1-\mathrm{P}(\mathrm{A} \cup \mathrm{B})}{\mathrm{P}\left(\mathrm{B}^{\prime}\right)}
    (D)
    P(A′)∣P(B′)\displaystyle P\left(A^{\prime}\right) \mid P\left(B^{\prime}\right)

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    C
    (C) \(\displaystyle \dfrac{1-P(A\cup B)}{P(B')}\)\[A'\cap B'=(A\cup B)' \quad \text{(De Morgan)} \] \[P(A'\mid B')=\frac{P(A'\cap B')}{P(B')}=\frac{P\big((A\cup B)'\big)}{P(B')}=\frac{1-P(A\cup B)}{P(B')} \]
  9. Exercise 69

    If A and B are two independent events with P(A)=35\displaystyle \mathrm{P}(\mathrm{A})=\frac{3}{5} and P(B)=49\displaystyle \mathrm{P}(\mathrm{B})=\frac{4}{9}, then P(A′∩B′)\displaystyle P\left(A^{\prime} \cap B^{\prime}\right) equals
    (A)
    415\displaystyle \frac{4}{15}
    (B)
    845\displaystyle \frac{8}{45}
    (C)
    13\displaystyle \frac{1}{3}
    (D)
    29\displaystyle \frac{2}{9}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    D
    (D) \(\displaystyle \dfrac{2}{9}\)A and B independent, so A' and B' are independent. \[P(A')=1-\frac35=\frac25,\qquad P(B')=1-\frac49=\frac59 \] \[P(A'\cap B')=P(A')\,P(B')=\frac25\cdot\frac59=\frac29 \]
  10. Exercise 70

    If two events are independent, then
    (A)
    they must be mutually exclusive
    (B)
    the sum of their probabilities must be equal to 1\displaystyle 1
    (C) (A) and (B) both are correct
    (D)
    None of the above is correct

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    D
    (D) None of the above is correct.Example: a coin shows head (A), a die shows six (B). \[P(A)=\frac12,\quad P(B)=\frac16,\quad P(A\cap B)=\frac1{12}=P(A)\,P(B) \] Independent, yet \[P(A\cap B)=\frac1{12}\neq 0 \quad\Rightarrow\quad \text{not mutually exclusive} \] \[P(A)+P(B)=\frac23\neq 1 \]